1.) The sum of the digits of a two digit no. is 12. The value of the number is five more than 10 times the tens digit. Find the number.
Answer by stanbon(75887)

The sum of the digits of a two digit no. is 12. The value of the number is five more than 10 times the tens digit. Find the number.
----------------
Let the number be 10t+u
------
Equations:
t + u = 12
10t+u = 10t+5
----
Solve the bottom equation::
u = 5
Then t = 12-5 = 7
----
Ans: 10t+u = 75
--------------------------
Cheers,
Stan H.
-------------------